OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (anonymous):

Number of sides \(= \frac {360} {72} = 5\)

OpenStudy (anonymous):

Now, use \((2n-4)\times 90 = 540^\circ , n = 5\)

OpenStudy (lgbasallote):

just wondering where does \[\large \frac{360}{72}\] come from @FoolForMath a formula?

OpenStudy (anonymous):

Sum of all exterior angles of any polygon is \(360^\circ \). In regular they are equal.

OpenStudy (lgbasallote):

i see thanks